. A long spring is stretched by 2cm. Its potential energy is V. If the spring is stretched by 10cm, its potential energy would be A. V/25 B. V/5 C. 5V D. 25V

Grade:upto college level

1 Answers

Unni
5 years ago
potential energy of spring =1/2kx^2
Therefore when x=2 ,PE=1/2k(2^2)=2k
Similarly x=10,PE=1/2k(10^2)=50k=25×2k
Given 2k=U then 25×2k=25U (I given v=u)